# Energy Conditions and Quantum Information

### v1

@@ 3 Energy conditions and quantum field theory
In this section, we turn to energy conditions in quantum field theory in Minkowski spacetime. After giving a brief review of quantum information in a bipartite system, we consider the modular Hamiltonian $K_{A}$ associated with the subsystem $A$ in the vacuum. When the boundary of $A$ is planar, the Bisognano-Wichmann theorem Bisognano:1976za states that $K_{A}$ generates the Rindler boost, hence be written by smearing the stress-energy tensor on $A$. This theorem generalizes to the case when $A$ has the boundary on a null hypersurface Casini:2017roe , which allows us to derive the ANEC from the monotonicity of the vacuum modular Hamiltonian Faulkner:2016mzt . We also consider the quantum null energy condition (QNEC), which bounds the null-null component of the stress tensor by the second null derivative of entanglement entropy Bousso:2015mna ; Wall:2011kb . We sketch the derivation of QNEC following Kudler-Flam:2023hkl and leave the complete proof to Ceyhan:2018zfg ; Hollands:2025glm .
@@ 3.1 Quantum information measures
$S(\rho_{A}||\sigma_{A})\equiv{\rm tr}_{A}\left[\rho_{A}\,\log\rho_{A}\right]-{\rm tr}_{A}\left[\rho_{A}\,\log\sigma_{A}\right]\ .$ (58)
@@
In fact, this expression coincides with the free energy $F=\Delta E-T\Delta S$, if $\sigma_{A}$ is the thermal density matrix.
When the system $\bar{A}$ is a purifying system of both $\rho_{A}$ and $\sigma_{A}$, we can simplify $\Delta K^{(\sigma)}_{A}$ to the following form:
@@ 3.2 Modular Hamiltonian in quantum field theory
The modular Hamiltonian defined by (62) does not have a local expression written by an integral of a local operator in general, unlike a Hamiltonian. However, when $\sigma$ is the vacuum density matrix $\Omega$, i.e.  $|\sigma=\Omega=\ket{\Omega}\bra{\Omega}$, the modular Hamiltonian (62) can be “geometric”, which is a local integral of the stress tensor, generating the time evolution. A well-known example is the Bisognano-Wichmann theorem Bisognano:1976za . Let region $A$ be the right half-space $A=\{t=0,x_{1}\geq 0\}$. We take the coordinate in Minkowski spacetime as
$t=r\,\sinh\tau\ ,\qquad x_{1}=r\,\cosh\tau\ .$ (66)
@@
That is, the coordinate $(r,\tau)$ implies the Minkowski coordinate of the observer who is uniformly accelerating in the Minkowski spacetime, and covers only $|t|\leq x_{1}$, known as the (right) Rindler wedge (see the left panel of figure 12). The Rindler wedge is the causal domain (diamond) $D[A]$ of the region $A$. The causal domain $D[A]$ is the set of points $p$ such that all causal curves through $p$ intersect $A$, and the information at the point $p$ is completely determined by that of $A$. The Hamiltonian $K$ for the observer with respect to the Rindler time $\tau$ is given by
$K=\int_{A=\{t=0,\,x_{1}\geq 0\}}{\rm d}^{d-1}x\,x_{1}\,T_{tt}\ .$ (67)
The Bisognano-Wichmann theorem states that the vacuum modular Hamiltonian is the Hamiltonian for the uniformly accelerating observer, i.e. $K_{A}^{(\Omega)}=2\pi\,K$. The physical interpretation of this theorem is as follows. Starting from $x_{1}\geq 0$ at the initial time, the uniformly accelerating observer cannot obtain the information about $x_{1}\leq 0$ at the same time, since even the signal with the speed of light never reaches the observer (for this observer, $x_{1}<0$ is regarded as the inside of the black hole). Thus, the system with this observer is described by the vacuum reduced density matrix $\sigma_{A}$, even if the total system is in a pure state (vacuum). Furthermore, $\Omega_{A}=e^{-2\pi K}$ implies that the Rindler observer perceives a thermal bath at inverse temperature $\beta=2\pi$.
@@
Figure 12: [Left] The right Rindler wedge is the causal domain $D[A]$ of the region $A$ as shown in gray color. The vacuum modular Hamiltonian coincides with the boost operator $K_{A}^{(\Omega)}=2\pi\,K$ due to the Bisognano-Wichmann theorem. [Right] The right Rindler wedge is mapped to the causal domain of the spherical region by the conformal transformation (68).
@@
While the Bisognano-Wichmann theorem holds in any relativistic quantum field theory, there is another expression for the modular Hamiltonian in conformal field theory (CFT). Let region $A$ be the solid ball of radius $R$ centered at the origin: $A=\{X^{0}=0,\,0\leq r\leq R\}$. Here, $r\equiv\sqrt{\sum_{i=1}^{d-1}(X^{i})^{2}}$ is the radius of the spherical coordinates in the constant-time slice. The Rindler wedge is mapped to the causal domain $D[A]$ of the solid ball $A$, by conformal transformation, called the CHM map Casini:2011kv (figure 12):
@@
where $c^{\mu}=(0,1/(2R),\vec{0}\,)$. This means that, in CFT, the vacuum reduced density matrix of the right half-space is related to $\Omega_{A}$ by a unitary transformation. Then, the modular Hamiltonian of $\Omega_{A}$ is given by Casini:2011kv ; Hislop:1981uh
$K^{(\Omega)}_{A}=2\pi\int_{X^{0}=0,\,0\leq r\leq R}{\rm d}^{d-1}X\,\frac{R^{2}-r^{2}}{2R}\,T_{00}\ .$ (69)
We can consider the null deformed Rindler wedge as the causal domain $D[A]$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ as shown in figure 13. The modular Hamiltonian associated with such a region $A$ is also shown in Casini:2017roe to have a local expression similar to (67)
Figure 13: The null deformed Rindler wedge as the causal domain $D[A]$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ .
Having the applications to QFTs in mind, it is useful to extend the monotonicity of relative entropy (61) to the operator inequality Borchers:1995zg ; Blanco:2013lea :
$D[B]\subset D[A]\quad\rightarrow\quad\hat{K}^{(\rho)}_{B}\leq\hat{K}^{(\rho)}_{A}\$, (71)
where $D[A]$ is the causal domain of a spacelike hypersurface $A$, and $\hat{K}^{(\rho)}_{A}$ is the full modular Hamiltonian
$\hat{K}^{(\rho)}_{A}$ $\equiv K^{(\rho)}_{A}\otimes{\bf 1}_{\bar{A}}-{\bf 1}_{A}\otimes K^{(\rho)}_{\bar{A}}\$. (72)
@@ 3.3 Bekenstein bound
where $S$ and $E$ are the entropy of the matter and energy in the region $A$ with typical size $R$. This bound arose through the following thought experiment. Consider an object with energy $E$ and entropy $S$, and drop into the black hole of radius $R$ in the three-dimensional space for simplicity (a similar argument holds in an arbitrary dimension). It is known that the black hole entropy is proportional to its area. As the radius of the black hole increases by $\Delta R\propto E$, the entropy also increases by $\Delta S_{\text{BH}}\propto R\Delta R$. The second law of thermodynamics implies that the entropy of the total system including an object and the black hole is bounded: $S\leq\Delta S_{\text{BH}}\propto R\,E$.
@@
To give a well-defined formulation of the Bekenstein bound, we apply the Bisognano-Wichmann theorem Casini:2008cr . Suppose that the region $A$ is the solid ball of the radius $R$ in CFT. Consider the reduced density matrices of an excited state and the vacuum state: $\rho_{A},\Omega_{A}$. The positivity of the relative entropy (59) and the relation to modular Hamiltonian (63) show
@@
where the $K^{(\Omega)}_{A}$ in the right hand side is the vacuum modular Hamiltonian in CFT (69). Let $E$ be the energy of the excited state in the region $A$, then we can estimate $T_{tt}\sim E/R^{d-1}$, since $T_{tt}$ is the energy density. Integrating over the region $A$, we obtain $\Delta K_{A}^{(\Omega)}\sim R\,E$. Therefore, we can think of (74) as the quantum version of the Bekenstein bound (73). In addition, a different form of the Bekenstein bound was discovered using the monotonicity of the relative entropy in Blanco:2013lea .
@@ 3.4 Averaged null energy condition
where we use the light-cone coordinates $x^{\mu}=(x^{+},x^{-},\bm{x}_{\perp})\ ,~x^{\pm}\equiv t\pm x_{1}$ in Minkowski spacetime (see also Definition 2.7.2). The ANEC holds for any QFT in Minkowski spacetime, which was proved in special cases (free theory, two dimensions, holography) Klinkhammer:1991ki ; Wald:1991xn ; Folacci:1992xg ; Verch:1999nt ; Kelly:2014mra , and in interacting theories Faulkner:2016mzt ; Hartman:2016lgu . Here, we give a proof using the monotonicity of the relative entropy along the line of Faulkner:2016mzt .
The modular Hamiltonian (70) for the null deformed Rindler wedge can also be written as
$K^{(\Omega)}_{A}=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{0}^{\infty}{\rm d}x^{+}x^{+}\,T_{++}\left(x^{+}-\gamma_{A}(\bm{x}_{\perp}),\,x^{-}=0,\,\bm{x}_{\perp}\right)\ .$ (76)
We also consider another region $B$, whose boundary satisfies $\partial B=\{x^{-}=0,x^{+}=\gamma_{B}(\bm{x}_{\perp})\}$ and $\Delta\gamma(\bm{x}_{\perp})\equiv\gamma_{B}(\bm{x}_{\perp})-\gamma_{A}(\bm{x}_{\perp})>0$ (figure 14). Then, since the causal diamonds have a relation $D[B]\subset D[A]$, the monotonicity of the full modular Hamiltonian (71) yields the inequality
@@ 3.5 Quantum null energy condition
While ANEC is a kind of global energy conditions, there is a local bound on the null energy density. This is called the quantum null energy condition (QNEC) Bousso:2015mna , whose statement is as follows: {itembox}[l]Definition 3.5.1 (QNEC)
@@
Here $|]\rho_{A}={\rm tr}_{\bar{A}}\left[\,\ket{\rho}\bra{\rho}\,\right$] and $S(\rho_{A})$ is the entanglement entropy of a spacelike hypersurface $A$ in $x^{-}\leq 0$ whose boundary is anchored on $x^{-}=0,x^{+}=\gamma_{A}(\bm{x}_{\perp})$. Note that the second-order null shape derivative of the entanglement entropy in the right-hand side is not necessarily positive. ANEC follows from QNEC by integrating both sides of (78) along, assuming that the surface term $\delta S/\delta\gamma$ vanishes. QNEC arose in the context of quantum focusing conjecture (QFC) Bousso:2015mna , and was proven in free theory Bousso:2015wca ; Malik:2019dpg , in holographic theories Koeller:2015qmn , and in interacting theories Balakrishnan:2017bjg ; Ceyhan:2018zfg ; Hollands:2025glm . [^12] For interacting CFT, QNEC is conjectured to be saturated Leichenauer:2018obf ; Balakrishnan:2019gxl . QNEC also has applications to quenches Mezei:2019sla ; Kibe:2024icu ; Kibe:2025cqc , and renormalization group flow Casini:2023kyj .
Let us give a sketch of a proof of QNEC (our proof mainly follows that of Kudler-Flam:2023hkl ). We will use the expression of the relative entropy (63) and the vacuum modular Hamiltonian (70). By taking $\sigma$ in (63) to be the vacuum reduced density matrix, the second-order null shape derivative of the relative entropy becomes
@@
This positivity is guaranteed by the “ant-formula”:

### v2

@@ 3 Energy conditions and quantum field theory
In this section, we turn to energy conditions in quantum field theory in Minkowski spacetime. After giving a brief review of quantum information in a bipartite system, we consider the modular Hamiltonian $K_{A}$ associated with the subsystem $A$ in the vacuum. When the boundary of $A$ is planar, the Bisognano-Wichmann theorem Bisognano:1976za states that $K_{A}$ generates the Rindler boost, hence be written by smearing the stress-energy tensor on $A$. This theorem generalizes to the case when $A$ has the boundary on a null hypersurface Casini:2017roe , which allows us to derive the ANEC from the monotonicity of the vacuum modular Hamiltonian Faulkner:2016mzt . We also consider QNEC, which bounds the null-null component of the stress tensor by the second null derivative of entanglement entropy Bousso:2015mna ; Wall:2011kb . We sketch the derivation of QNEC following Kudler-Flam:2023hkl and leave the complete proof to Ceyhan:2018zfg ; Hollands:2025glm .
@@ 3.1 Quantum information measures
$S(\rho_{A}||\sigma_{A}):={\rm tr}_{A}\left[\rho_{A}\,\log\rho_{A}\right]-{\rm tr}_{A}\left[\rho_{A}\,\log\sigma_{A}\right]\ .$ (58)
@@
In fact, this expression coincides with the free energy $F=\Delta E-T\Delta S$, if $\sigma_{A}$ is the thermal density matrix. When the system $\bar{A}$ is a purifying system of both $\rho_{A}$ and $\sigma_{A}$, we can simplify $\Delta K^{(\sigma)}_{A}$ to the following form:
@@
Having applications to QFTs in mind, we introduce the full modular Hamiltonian $\hat{K}^{(\rho)}_{A}$ as follows:
$\hat{K}^{(\sigma)}_{A}$ $:=K^{(\sigma)}_{A}\otimes{\bf 1}_{\bar{A}}-{\bf 1}_{A}\otimes K^{(\sigma)}_{\bar{A}}\$. (66)
The monotonicity of relative entropy (61) can be expressed as the operator inequality for the full modular Hamiltonian Borchers:1995zg ; Blanco:2013lea :
$B\subset A\quad\rightarrow\quad\hat{K}^{(\sigma)}_{B}\leq\hat{K}^{(\sigma)}_{A}\$. (67)
@@ 3.2 Modular Hamiltonian in quantum field theory
From now on, we will extend the discussion in the previous section to quantum field theory. Let the Hilbert space $\mathcal{H}$ be defined on a Cauchy surface $\Sigma$, and denote by $\mathcal{H}_{A}$ the Hilbert space associated with a subregion $A\subset\Sigma$. In a finite-dimensional system, the Hilbert space factorizes as $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}$, which allows us to define the reduced density matrix $\rho_{A}$, modular Hamiltonian $K_{A}^{(\sigma)}$ and entanglement entropy. In quantum field theory, however, such a factorization no longer holds, so these quantities are not well-defined in general. Nevertheless, in an algebraic formulation of QFT, one can rigorously define the relative entropy $S(\rho_{A}||\sigma_{A})$ and the full modular Hamiltonian $\hat{K}_{A}^{(\sigma)}$, which generates the modular flow of local operator algebra (see witten2018notes-8d6 for a review).
In what follows, we will take a less rigorous approach and assume that the Hilbert space factorizes under a suitable ultraviolet regularization, so that the reduced density matrix, modular Hamiltonian, and entanglement entropy can be treated for practical purposes.
@@
The full modular Hamiltonian defined by (66) does not have a local expression written by an integral of a local operator in general, unlike a Hamiltonian. However, when $\sigma$ is the vacuum density matrix $\Omega$, i.e.  $|\sigma=\Omega=\ket{\Omega}\bra{\Omega}$, the full modular Hamiltonian (66) can be “geometric,” which is a local integral of the stress tensor, generating the time evolution. A well-known example is the Bisognano-Wichmann theorem Bisognano:1976za . Let region $A$ be the right half-space $A=\{t=0,x^{1}\geq 0\}$. We take the coordinate in Minkowski spacetime as
$t=r\,\sinh\tau\ ,\qquad x^{1}=r\,\cosh\tau\ .$ (68)
@@
That is, the coordinate $(r,\tau)$ implies the Minkowski coordinate of the observer who is uniformly accelerating in the Minkowski spacetime, and covers only $|t|\leq x^{1}$, known as the (right) Rindler wedge (see the left panel of figure 12). The Rindler wedge is the causal domain (diamond) $D(A)$ of the region $A$. The causal domain $D(A)$ is the set of points $p$ such that all causal curves through $p$ intersect $A$, and the information at the point $p$ is completely determined by that of $A$. The Hamiltonian $K$ with respect to the Rindler time $\tau$ is given by
$K=\int_{t=0}{\rm d}^{d-1}x\,x^{1}\,T_{tt}\ .$ (69)
[…]
$K_{R}=\int_{A=\{t=0,\,x^{1}\geq 0\}}{\rm d}^{d-1}x\,x^{1}\,T_{tt}\ .$ (70)
On the other hand, assuming that quantum field theory has a tensor factorized Hilbert space under some UV regularization, the system of the observer may be characterized by the reduced density matrix $\Omega_{A}=e^{-K_{A}^{(\Omega)}}$, where $K_{A}^{(\Omega)}$ is the modular Hamiltonian. The Bisognano-Wichmann theorem then implies $K_{A}^{(\Omega)}=2\pi\,K_{R}$, so that $\Omega_{A}=e^{-2\pi K_{R}}$. This shows that the Rindler observer perceives a thermal bath at inverse temperature $\beta=2\pi$, even though the total system is in a pure state.
@@
Figure 12: [Left] The right Rindler wedge is the causal domain $D(A)$ of the region $A$ as shown in gray color. The vacuum modular Hamiltonian coincides with the boost operator $K_{A}^{(\Omega)}=2\pi\,K_{R}$ due to the Bisognano-Wichmann theorem. [Right] The right Rindler wedge is mapped to the causal domain of the spherical region by the conformal transformation (71).
@@
While the Bisognano-Wichmann theorem holds in any relativistic quantum field theory (under certain condition), there is another expression for the modular Hamiltonian in conformal field theory (CFT). Let region $A$ be the solid ball of radius $R$ centered at the origin: $A=\{X^{0}=0,\,0\leq r\leq R\}$. Here, $r:=\sqrt{\sum_{i=1}^{d-1}(X^{i})^{2}}$ is the radius of the spherical coordinates in the constant-time slice. The Rindler wedge is mapped to the causal domain $D(A)$ of the solid ball $A$, by conformal transformation, called the CHM map Casini:2011kv (figure 12):
@@
where $x^{0}=t$ and $c^{\mu}=(0,1/(2R),\vec{0}\,)$. This means that, in CFT, the vacuum reduced density matrix of the right half-space is related to $\Omega_{A}$ by a unitary transformation. Then, the modular Hamiltonian of $\Omega_{A}$ is given by Casini:2011kv ; Hislop:1981uh
$K^{(\Omega)}_{A}=2\pi\int_{X^{0}=0,\,0\leq r\leq R}{\rm d}^{d-1}X\,\frac{R^{2}-r^{2}}{2R}\,T_{X^{0}X^{0}}\ .$ (72)
It follows that the full modular Hamiltonian $\hat{K}^{(\Omega)}_{A}$ takes the same form as (72) with the integration over the entire time slice $X^{0}=0$. While the derivation of $\hat{K}^{(\Omega)}_{A}$ was not fully rigorous, the resulting expression can be shown to hold exactly in the case of a free massless scalar field Longo:2020amm .
Let us define the light-cone coordinates $x^{\mu}=(x^{+},x^{-},\bm{x}_{\perp})\ ,~x^{\pm}:=t\pm x^{1}$ in Minkowski spacetime. We can consider the null deformed Rindler wedge as the causal domain $D(A)$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ as shown in figure 13. The full modular Hamiltonian associated with such a region $A$ in quantum field theory is also shown, at a physicist’s level of rigor in Casini:2017roe , to have a local expression similar to (70):
$\hat{K}^{(\Omega)}_{A}=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{-\infty}^{\infty}{\rm d}x^{+}(x^{+}-\gamma_{A}(\bm{x}_{\perp}))\,T_{++}\left(x^{+},x^{-}=0,\bm{x}_{\perp}\right)\ .$ (73)
We will use this expression in deriving averaged null energy condition on flat space in section 3.4.
Figure 13: The null deformed Rindler wedge as the causal domain $D(A)$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ .
@@ 3.3 Bekenstein bound
where $S$ and $E$ are the entropy of the matter and energy in the region $A$ with typycal size $R$. This bound arose through the following thought experiment. Consider an object with energy $E$ and entropy $S$, and drop into the black hole of radius $R$ in the three-dimensional space for simplicity (a similar argument holds in an arbitrary dimension). It is known that the black hole entropy is proportional to its area. As the radius of the black hole increases by $\Delta R\propto E$, the entropy also increases by $\Delta S_{\text{BH}}\propto R\Delta R$. The second law of thermodynamics implies that the entropy of the total system including an object and the black hole is bounded: $S\leq\Delta S_{\text{BH}}\propto R\,E$.
@@
To give a well-defined formulation of the Bekenstein bound, we apply the Bisognano-Wichmann theorem Casini:2008cr . Suppose that the region $A$ is the solid ball of the radius $R$ in CFT. Consider the reduced density matrices of an excited state and the vacuum state: $\rho_{A},\Omega_{A}$. The positivity of the relative entropy (59) and the relation to the modular Hamiltonian (63) show
@@
where the $K^{(\Omega)}_{A}$ in the right hand side is the vacuum modular Hamiltonian in CFT (72). Let $E$ be the energy of the excited state in the region $A$, then we can estimate $T_{X^{0}X^{0}}\sim E/R^{d-1}$, since $T_{X^{0}X^{0}}$ is the energy density. Integrating over the region $A$, we obtain $\Delta K_{A}^{(\Omega)}\sim R\,E$. Therefore, we can think of (75) as the quantum version of the Bekenstein bound (74). In addition, a different form of the Bekenstein bound was discovered using the monotonicity of the relative entropy in Blanco:2013lea .
@@ 3.4 Averaged null energy condition
The ANEC holds for any QFT in Minkowski spacetime, which was proved in special cases (free theory, two dimensions, holography) Klinkhammer:1991ki ; Wald:1991xn ; Folacci:1992xg ; Verch:1999nt ; Kelly:2014mra , and in interacting theories Faulkner:2016mzt ; Hartman:2016lgu . Here, we give a proof using the monotonicity of the relative entropy along the line of Faulkner:2016mzt .
We begin with the full modular Hamiltonian (73) for the null deformed Rindler wedge $D(A)$. We also consider another region $B$, whose boundary satisfies $\partial B=\{x^{-}=0,x^{+}=\gamma_{B}(\bm{x}_{\perp})\}$ and $\Delta\gamma(\bm{x}_{\perp}):=\gamma_{B}(\bm{x}_{\perp})-\gamma_{A}(\bm{x}_{\perp})>0$ (figure 14). In quantum field theory, (67) implies
$D(B)\subset D(A)\quad\rightarrow\quad\hat{K}^{(\rho)}_{B}\leq\hat{K}^{(\rho)}_{A}\$, (77)
where $D(A)$ and $D(B)$ are the causal domains of spacelike hypersurfaces $A$ and $B$, respectively. Thus, in the present setup, (73) and (77) yield the inequality
@@ 3.5 Quantum null energy condition
While ANEC is a kind of global energy conditions, there is a local bound on the null energy density. This is called the quantum null energy condition (QNEC) Bousso:2015mna , whose statement is as follows:
@@
Let us give a sketch of a physicist’s derivation of QNEC (we closely follows that of Kudler-Flam:2023hkl ). We will use the expression of the relative entropy (63) and the vacuum modular Hamiltonian corresponding to (73)
By taking $\sigma$ in (63) to be the vacuum reduced density matrix, the second-order null shape derivative of the relative entropy becomes
@@
This inequality is mathematically well-defined while the original QNEC (79) is not as the latter involves entanglement entropy explicitly. (82) is guaranteed by the “ant-formula”: